# HopaDIFF: Holistic-Partial Aware Fourier Conditioned Diffusion for Referring Human Action Segmentation in Multi-Person Scenarios

### v1

@@ 5.3 Ablation Analysis
The variants include: (1) w/o all, a vanilla diffusion-based action segmentation model without any of the proposed components; (2) w/o HCMGB, which removes the GroundingDINO [34]-based partial reasoning branch; (3) w/o HP-xLSTM, which omits the holistic-partial aware xLSTM temporal aggregation module; (4) w/o BCAF, which excludes the bidirectional cross-attention fusion between the input gates of the two branches; and (5) w/o DFT-cond, which eliminates frequency-domain conditioning in both branches. The full HopaDIFF model achieves the highest performance, with $59.63\%$ ACC, $19.37\%$ EDIT, and $90.91\%$, $90.33\%$, $89.26\%$ F1 scores at thresholds $10\%$, $25\%$, and $50\%$ on the test set. Removing HP-xLSTM leads to a significant drop in F1@50 from $89.26\%$ to $83.50\%$, demonstrating the importance of long-range holistic-partial temporal modeling. Eliminating DFT-cond lowers F1@50 to $86.38\%$, validating the effectiveness of frequency-based conditioning in improving the precision and controllability of action segmentation generation. The absence of BCAF also causes a performance decrease, with F1@50 dropping to $81.38\%$, highlighting the utility of cross-branch interaction through input gate attention. Notably, removing the GroundingDINO [34]-based partial reasoning branch (HCMGB) reduces F1@50 to $49.42\%$, showing that person-specific local cues are essential for accurate target-aware segmentation. Overall, the ablation results confirm that each proposed component, i.e., GroundingDINO [34]-based partial cue reasoning, HP-xLSTM-based fusion, and frequency conditioning, plays a critical role in our model, and their synergy leads to the superior performance of HopaDIFF on the complex textual referring action segmentation task. More ablations can be found in the supplementary.
@@ 6 Conclusion
In this work, we introduced Referring Human Action Segmentation (RHAS), a new task aimed at segmenting actions of individuals in multi-person untrimmed videos using natural language references. To support this, we constructed the RHAS133 dataset with fine-grained action labels and referring expressions. We proposed HopaDIFF, a dual-branch diffusion model that fuses global and localized cues via HP-xLSTM and DFT-based conditioning for improved temporal reasoning and control. Extensive experiments show that HopaDIFF outperforms existing baselines in both accuracy and temporal consistency. Our contributions pave the way for fine-grained, language-guided video understanding in complex settings. Limitations: The experiments are conducted on one dataset, i.e., RHAS, since there is no other textual reference human action segmentation dataset available in the community. In the future, we plan to deliver more datasets within this research direction to enable more generalizable evaluations.
@@ Acknowledgements
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation) with a start date of April 1, 2024. This work was also partially supported by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This project is also supported by the National Natural Science Foundation of China under Grant No. 62473139. Lastly, the authors thank for the support of Dr. Sepideh Pashami, the Swedish Innovation Agency VINNOVA, the Digital Futures.
@@ A.1.1 Limitations
As stated in our main paper, our work is the first to investigate human action segmentation guided by textual references to enable action segmentation for a specific individual in multi-person scenarios. Due to the novelty of the task, experiments are currently limited to our constructed dataset, which constrains the evaluation of the proposed model’s generalizability. Nevertheless, we made substantial efforts to ensure dataset diversity, collecting data from a wide range of movies over an annotation process that spanned more than $8$ months. To further assess generalizability, we designed multiple evaluation protocols, e.g., random partition and cross-movie partition, and conducted cross-backbone experiments by replacing BLIP-2 [26] with CLIP. In future work, we aim to expand the dataset by incorporating a larger and more diverse set of movies.
Appendix B More Implementation Details
In this section, we will provide more implementation details regarding our proposed HopaDIFF framework. For both encoders of the holistic branch and the partial branch, the number of feature maps is chosen as $64$, and the kernel size is chosen as $5$. We adopt a normal dropout rate of $0.5$, a channel dropout rate of $0.5$, and a temporal dropout rate of $0.5$. For both the decoders of the holistic branch and the partial branch, we choose the dimension of time embedding as $512$, the number of feature maps as $24$, the kernel size as $5$, and the dropout rate as $0.1$. Regarding the hyperparameters of the diffusion process, we choose the timesteps as $1,000$ and the sampling time steps as $25$. The training epoch is set as $1,000$.
Appendix C Ablation Study of the frame length
@@
Appendix D More Details of RHAS133 Dataset
In this section, we deliver more details of our contributed RHAS133 dataset. First, we present the number of persons statistic for each movie in Fig. 4, which illustrates that our dataset contains diverse distribution regarding the number of persons per video.

### v2

@@ 5.3 Ablation Analysis
Table 6: Ablation experiments for partial branch, when we use BLIP-2 and cross-movie partition.
Method Val Test
ACC EDIT F1@10 F1@25 F1@50 ACC EDIT F1@10 F1@25 F1@50
FACT lu2024fact 16.73 0.39 66.07 65.73 64.69 15.82 0.30 57.30 55.75 53.72
ActDIFF liu2023diffusion 3.84 3.51 51.38 50.53 47.87 2.04 7.67 44.13 43.70 42.42
ASQuery gan2024asquery 32.42 0.11 51.18 48.31 43.32 36.51 0.12 44.46 42.52 39.09
LTContent bahrami2023much 45.36 0.08 60.51 57.00 50.87 50.34 0.08 49.58 47.14 42.43
RefAtomNet peng2024referring 27.73 0.13 46.90 43.52 38.82 30.45 0.15 38.01 36.30 32.72
Ours w/ only PB 44.48 6.28 79.40 77.34 74.42 48.45 19.70 81.99 81.49 80.48
Ours w/ HPB 54.41 6.31 89.14 86.56 83.35 59.63 19.37 90.11 90.33 89.26
The ablated module variants include: (1) w/o all, a vanilla diffusion-based action segmentation model without any of the proposed components; (2) w/o HCMGB, which removes the GroundingDINO liu2024grounding -based partial reasoning branch; (3) w/o HP-xLSTM, which omits the holistic-partial aware xLSTM temporal aggregation module; (4) w/o BCAF, which excludes the bidirectional cross-attention fusion between the input gates of the two branches; and (5) w/o DFT-cond, which eliminates frequency-domain conditioning in both branches. The full HopaDIFF model achieves the highest performance, with $59.63\%$ ACC, $19.37\%$ EDIT, and $90.91\%$, $90.33\%$, $89.26\%$ F1 scores at thresholds $10\%$, $25\%$, and $50\%$ on the test set. Removing HP-xLSTM leads to a significant drop in F1@50 from $89.26\%$ to $83.50\%$, demonstrating the importance of long-range holistic-partial temporal modeling. Eliminating DFT-cond lowers F1@50 to $86.38\%$, validating the effectiveness of frequency-based conditioning in improving the precision and controllability of action segmentation generation. The absence of BCAF also causes a performance decrease, with F1@50 dropping to $81.38\%$, highlighting the utility of cross-branch interaction through input gate attention. Notably, removing the GroundingDINO liu2024grounding -based partial reasoning branch (HCMGB) reduces F1@50 to $49.42\%$, showing that person-specific local cues are essential for accurate target-aware segmentation. Overall, the ablation results confirm that each proposed component, i.e., GroundingDINO liu2024grounding -based partial cue reasoning, HP-xLSTM-based fusion, and frequency conditioning, plays a critical role in our model, and their synergy leads to the superior performance of HopaDIFF on the complex textual referring action segmentation task. The experiments also reveal that replacing the action branch with the partial branch in FACT ((6) FACT lu2024fact w/ partial branch) improves F1@{10,25,50} by better capturing long-duration, movement-related actions, but decreases ACC due to the loss of broader contextual cues. Adding a frequency-domain condition to ActDIFF ((7) ActDIFF liu2023diffusion w/ DFT-cond) enhances boundary detection (higher EDIT) but shows weaker generalization across movies without the HP-xLSTM module, leading to lower F1 on validation. Finally, while cross-attention fusion performs competitively, our HP-xLSTM achieves superior accuracy and F1 by enabling structured, fine-grained temporal modeling and selective bidirectional information exchange between holistic and partial branches. We further test cropping the person region detected by GroundingDino before feeding it into the baselines (Table 6). Compared to holistic inputs (Table 3), this results in lower ACC but partial gains in F1, showing the complementary roles of holistic and partial branches: the holistic branch captures background and interaction cues, while the partial branch focuses on the target individual to improve movement-related action recognition.
@@ 6 Conclusion
In this work, we introduced Referring Human Action Segmentation (RHAS), a new task aimed at segmenting actions of individuals in multi-person untrimmed videos using natural language references. To support this, we constructed the RHAS133 dataset with fine-grained action labels and referring expressions. We proposed HopaDIFF, a dual-branch diffusion model that fuses global and localized cues via HP-xLSTM and DFT-based conditioning for improved temporal reasoning and control. Extensive experiments show that HopaDIFF outperforms existing baselines in both accuracy and temporal consistency. Our contributions pave the way for fine-grained, language-guided video understanding in complex settings.
@@ Acknowledgements
The project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1574 – 471687386. This work was also partially supported by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This project is also supported in part by the National Natural Science Foundation of China under Grant No. 62473139, in part by the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20). This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure).
Appendix
@@ A.1.1 Limitations
As stated in our main paper, our work is the first to investigate human action segmentation guided by textual references to enable action segmentation for a specific individual in multi-person scenarios. Due to the novelty of the task, experiments are currently limited to our constructed dataset, which constrains the evaluation of the proposed model’s generalizability. Nevertheless, we made substantial efforts to ensure dataset diversity, collecting data from a wide range of movies over an annotation process that spanned more than $8$ months. To further assess generalizability, we designed multiple evaluation protocols, e.g., random partition and cross-movie partition, and conducted cross-backbone experiments by replacing BLIP-2 [29] with CLIP. In future work, we aim to expand the dataset by incorporating a larger and more diverse set of movies. Our method is a two-stage method that relies on pre-extracted VLM features, limiting the efficiency of inference. However, we also note that most existing methods (e.g., FACT, ActDIFF, LTContext, ASQuery) also adopt a two-stage processing pipeline due to the high computational cost of long video processing. We acknowledge this point also as an open challenge in action segmentation field.
Appendix B Further Clarification of the BCA Mechanism
The BCA module enables information exchange between holistic and partial branches by using cross-input-gate attention to modulate input information for each stream jointly, supporting richer temporal reasoning. HP-xLSTM leverages cross-input gated attention (i.e., BCA applied to input gates), enabling each branch to selectively incorporate temporal cues from the other based on feature-level gating. This design enhances temporal segmentation and boosts per-frame accuracy, particularly in challenging multi-person scenarios with fine-grained actions.
The attention weights and fused representations are computed as follows:
$\alpha_{t}^{H\rightarrow P}=\text{softmax}\left(\frac{(W_{Q}X^{H}_{t})(W_{K}X^{P}_{t})^{\top}}{\sqrt{d}}\right),$ (8)
$\tilde{X}^{P}_{t}=\alpha_{t}^{H\rightarrow P}\cdot(W_{V}X^{H}_{t}),$ (9)
$\alpha_{t}^{P\rightarrow H}=\text{softmax}\left(\frac{(W_{Q}X^{P}_{t})(W_{K}X^{H}_{t})^{\top}}{\sqrt{d}}\right),$ (10)
$\tilde{X}^{H}_{t}=\alpha_{t}^{P\rightarrow H}\cdot(W_{V}X^{P}_{t}),$ (11)
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ are learnable projection matrices for query, key, and value. $\tilde{X}^{P}_{t}$ and $\tilde{X}^{H}_{t}$ are the cross-attended inputs of the xLSTM input gates in the corresponding branches.
Appendix C More Implementation Details
In this section, we will provide more implementation details regarding our proposed HopaDIFF framework. For both encoders of the holistic branch and the partial branch, the number of feature maps is chosen as $64$, and the kernel size is chosen as $5$.
Table 7: Efficiency measurement when we leverage BLIP-2 feature extractor on RHAS133, cross-movie partition.
Method Trainable Param. Inference Time GFLOPS
FACT [40] 10.29M 1.98s 16.37G
ACTDIFF [34] 7.7M 2.28s 7.14G
ASQuery [18] 47.28M 1.24s 57.22G
LTContent [2] 6.12M 2.70s 20.02G
RefAtomNet [42] 106M 3.29s 273.53G
HopaDIFF 20M 2.81s 10.72G
We adopt a normal dropout rate of $0.5$, a channel dropout rate of $0.5$, and a temporal dropout rate of $0.5$. For both the decoders of the holistic branch and the partial branch, we choose the dimension of time embedding as $512$, the number of feature maps as $24$, the kernel size as $5$, and the dropout rate as $0.1$. Regarding the hyperparameters of the diffusion process, we choose the timesteps as $1,000$ and the sampling time steps as $25$. The training epoch is set as $1,000$. The number of parameters, inference time, and GFLOPS of the baseline approaches and our approach are reported in Table 7.
Appendix D Ablation Study of the frame length
@@
Appendix E Ablation regarding the Dual Head Setting
Training two branches separately leads to better performance compared with the mentioned counterpart, as shown in Table 9. It presents an ablation study comparing the Merged baseline with our dual-head design. On the validation set, our approach achieves higher accuracy ($54.41\%$ vs. $52.12\%$) and consistently outperforms across all F1 metrics, with improvements of $2.04\%$, $2.05\%$, and $1.86\%$ at F1@10, F1@25, and F1@50, respectively. On the test set, our method similarly shows superior accuracy ($59.63\%$ vs. $56.50\%$) and notable gains in F1 scores across all thresholds. Although the merged setting achieves slightly higher EDIT ($6.42$ vs. $6.31$ on validation, $19.89$ vs. $19.37$ on test), our dual head design provides a better balance, enhancing both per-frame accuracy and temporal overlap performance. These results highlight the advantage of explicitly modeling holistic and partial branches separately rather than merging them, as the dual head structure allows more fine-grained temporal reasoning and better recognition of actions across different contexts.
Table 9: Ablation study of the dual head setting.
Method Val Test
ACC EDIT F1@10 F1@25 F1@50 ACC EDIT F1@10 F1@25 F1@50
Merged 52.12 6.42 87.10 84.51 81.49 56.50 19.89 88.28 87.69 86.64
Ours 54.41 6.31 89.14 86.56 83.35 59.63 19.37 90.11 90.33 89.26
Appendix F Ablation regarding the Diffusion Action Segmentation Mechanism
To evaluate the contribution of the proposed diffusion mechanism in our human action segmentation framework, we conduct an ablation study by removing the diffusion module and comparing the results against our full model. The quantitative results are reported in Table 10, covering both validation and test sets across multiple metrics.
Without the diffusion mechanism, the model demonstrates a clear degradation in performance across all metrics. On the validation set, the accuracy drops from $54.41\%$ to $53.43\%$, while the EDIT score declines drastically from $6.31$ to $0.09$, indicating poor temporal alignment. Similarly, F1 scores decrease significantly, e.g., from $89.14\%$ to $67.35\%$ at F1@10 and from $83.35\%$ to $81.50\%$ at F1@50. A similar trend is observed on the test set, where the absence of diffusion reduces the EDIT score from $19.37$ to $0.10$, and segmental F1 scores experience large drops, such as from $90.33\%$ to $56.17\%$ at F1@25.
These results highlight the importance of the diffusion mechanism for our model. The large improvement in the EDIT score shows that diffusion plays a critical role in refining temporal boundaries and improving sequence coherence. Furthermore, the consistent gains across all F1 metrics confirm that diffusion enhances the model’s ability to capture fine-grained action transitions, ultimately leading to more robust and accurate action segmentation.
Table 10: Ablation study of the diffusion-based human action segmentation mechanism.
Method Val Test
ACC EDIT F1@10 F1@25 F1@50 ACC EDIT F1@10 F1@25 F1@50
w/o Diffusion 53.43 0.09 67.35 63.68 81.50 57.47 0.10 58.56 56.17 53.72
Ours 54.41 6.31 89.14 86.56 83.35 59.63 19.37 90.11 90.33 89.26
Appendix G Ablation regarding the Efficacy of Different Branches
To further investigate the contribution of each component in our framework, we conduct an ablation study on the partial and holistic branches. The results are summarized in Table 11, where we compare the performance of our full model against two variants: one without the partial branch and one without the holistic branch.
When the partial branch is removed, the performance drops drastically across all metrics. On the validation set, the accuracy decreases from $54.41\%$ to $20.78\%$, and F1 scores suffer severe degradation, e.g., from $89.14\%$ to $45.81\%$ at F1@10 and from $83.35\%$ to $41.95\%$ at F1@50. A similar trend is observed on the test set, where accuracy falls from $59.63\%$ to $23.44\%$, and F1@25 decreases sharply from $90.33\%$ to $50.14\%$. These results highlight the critical role of the partial branch in capturing fine-grained and localized action cues.
Removing the holistic branch also results in performance degradation, though the drop is less severe than removing the partial branch. On the validation set, accuracy decreases to $44.48\%$, while F1@10 and F1@25 decline to $79.40\%$ and $77.34\%$, respectively. On the test set, the accuracy reduces to $48.45\%$, with F1@25 dropping from $90.33\%$ to $81.49\%$. This indicates that the holistic branch contributes to capturing global temporal dependencies and overall action context, thereby improving segmentation coherence.
Overall, these ablations demonstrate that both the partial and holistic branches play complementary roles in our method. The partial branch is essential for modeling detailed action transitions, while the holistic branch ensures global consistency. Their combination enables our model to achieve state-of-the-art segmentation performance.
Appendix H Qualitative Results
We further provide a visualization of qualitative results of our approach and the FACT model based on one sample from the RHAS133 dataset in Figure 5. We observe that HopaDIFF consistently produces action segmentation outputs that more closely align with the ground truth labels. Notably, our method demonstrates superior performance in handling periodic action transitions along the temporal axis. This improvement can be attributed to the high granularity of the condition offered by the Fourier embeddings, especially on periodic temporal patterns, and the enhanced temporal reasoning capabilities of our proposed holistic-partial-aware cross-input-gate xLSTM architecture for different visual focuses. We also identified a performance drop in classes with limited training samples. In such long-tail scenarios, frames corresponding to rare classes are often misclassified as more frequent classes. We acknowledge this limitation and consider tackling the long-tail distribution challenge in referring human action segmentation as a promising and important direction for future research.
Table 11: Ablation study of different branches in our method.
Method Val Test
ACC EDIT F1@10 F1@25 F1@50 ACC EDIT F1@10 F1@25 F1@50
w/o partial branch 20.78 6.34 45.81 43.93 41.95 23.44 19.93 50.60 50.14 49.42
w/o holistic branch 44.48 6.28 79.40 77.34 74.42 48.45 19.70 81.99 81.49 80.48
Ours 54.41 6.31 89.14 86.56 83.35 59.63 19.37 90.11 90.33 89.26
@@ Appendix H Qualitative Results
Figure 5: Qualitative results of our HopaDIFF and FACT baseline, where false predictions are marked as green and correct predictions are marked as blue. Each color shown in GT denotes a different set of combinations of atomic-level actions.
Appendix I More Details of the RHAS133 Dataset
In this section, we deliver more details of our contributed RHAS133 dataset. First, we present the number of persons statistic for each movie in Fig. 6, which illustrates that the distribution of our dataset is diverse in terms of the number of persons per video.